Let
A=2x^2i−3yzj+xz^2k
and
ϕ=2z−x^3y
, find
A×▽ϕ
at point (1,-1,1).
1
Expert's answer
2021-09-30T14:20:08-0400
"\nabla \phi=-3x^2yi-x^3j+2k"

"A\times \nabla \phi=\begin{vmatrix}
 i & j & k \\
 2x^2 & -3yz & xz^2 \\
-3x^2y & -x^3 & 2
\end{vmatrix}"

"=i\begin{vmatrix}
 -3yz & xz^2 \\
 -x^3 & 2
\end{vmatrix}-j\begin{vmatrix}
 2x^2 & xz^2 \\
 -3x^2y & 2
\end{vmatrix}+k\begin{vmatrix}
 2x^2 & -3yz \\
 -3x^2y & -x^3
\end{vmatrix}"

"=(-6yz+x^4z^2)i+(-4x^2-3x^3yz^2)j"

"+(-2x^5-9x^2y^2z)k"

"(1,-1,1)"


"A\times \nabla \phi|_{(1, -1, 1)}=(6+1)i+(-4+3)j+(-2-9)k"

"=7i-j-11k"